OpenStudy (anonymous):

Malorie can work 10 math problems in 3 minutes. At this rate, how many problems can she work in one hour? A. 30 B. 60 C. 120 D. 200 Please help!

OpenStudy (anonymous):

\[ \frac{ 10 math problems }{ 3 minutes }\times \frac{ 60 \min }{ 1 hour }\]

jimthompson5910 (jim_thompson5910):

there are 60 minutes in an hour so the goal is to somehow get the "3 minutes" to turn into "60 minutes" how do we do this? we multiply "3 minutes" by 20 since 3*20 = 60 so we do the same to the other part as well ------------------------------------------------------- 10 math problems : 3 minutes 20*(10 math problems) : 20*(3 minutes) ... multiply both parts by 20 (20*10) math problems : (20*3) minutes 200 math problems : 60 minutes 200 math problems : 1 hour

OpenStudy (anonymous):

Oh my gosh, that's the best! Thank you! I suck at this math, and that is super helpful!!!!

jimthompson5910 (jim_thompson5910):

I'm glad it's making more sense now